OpenStudy (anonymous):

can someone check my answer, will fan + medal Using the given zero, find all other zeros of f(x). -2i is a zero of f(x) = x4 - 45x2 - 196 a 2i, 14i, -14i b 2i, 7, -7 c 2i, 14, -14 d 2i, 7i, -7i

OpenStudy (anonymous):

is it c or a?

OpenStudy (anonymous):

I can help you

OpenStudy (anonymous):

Thanks, is the answer c? because i know -14 x 14 = 196

OpenStudy (anonymous):

yes the answer is c

OpenStudy (anonymous):

can you help with one more please?

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

whatever you need I am here :)

OpenStudy (anonymous):

For the given function, find the vertical and horizontal asymptote(s) (if there are any). f(x) = 2x^2 +1 / x^2 - 4 I know the Horiziontal Asymptote is y= 2 but can you solve for the vertical bc i dont know how to do it exactly the answer choices: None x = 2, y = 2, y = 0 x = 2, x = -2, y = 2 x = 2, y = 2, y = 1

OpenStudy (anonymous):

to find vertical asymptotes you must find 0's in the denominator. so denominator (im guessing) is x^2-3x-4. so you factor out. getting (x-4)(x+1)=0 because we need one of those to be zero for an asymptote. x-4=0, x=4. x+1=0, x=-1. answer VA's occur when x= 4,-1. (ps. remember to check the numerator for the same zeros at the bottom. they cancel out and create holes not asymptotes.) But the answer is B